OpenStudy (anonymous):

A bag contains 2 steel balls and 5 brass balls. The total weight is 13 pounds. If 2 steel balls are added and 2 brass balls are removed, the weight decreases to 12 pounds. How much does each kind of ball weigh? A. steel, 4 pounds; brass, 1 poundB. steel, 1.5 pounds; brass, 2 poundsC. steel, 0.75 pounds; brass, 3 poundsD. steel, 3 pounds; brass, 4 pounds

OpenStudy (mathmate):

Long way: set up system of equations and solve 2S+5B=13 4S+3B=12 Solve for S and B.

OpenStudy (anonymous):

i already solved for it... i went for c

OpenStudy (mathmate):

Sorry, C is not correct.

OpenStudy (mathmate):

Try to put 2*0.75+5*3 to see if you get 13?

OpenStudy (mathmate):

Do you want to know the short way?

OpenStudy (anonymous):

What is the answer to this?

OpenStudy (anonymous):

@mathmate plea help?

OpenStudy (anonymous):

please* I want to learn the short way, lol

OpenStudy (anonymous):

@AccessDenied @Hero Please help me solve this one? :)

OpenStudy (anonymous):

its 1.5

OpenStudy (mathmate):

@tacosaur Better late than never! The short way for this: A bag contains 2 steel balls and 5 brass balls. The total weight is 13 pounds. If 2 steel balls are added and 2 brass balls are removed, the weight decreases to 12 pounds. How much does each kind of ball weigh? 2S+5B=13 we know that \(adding\) 2S-2B decreases by 1 lb. so \(subtracting\) 2S-2B will increase by 1 lb. Then we end up with 0S +7B = 13+1=14 lb, or B=2 lb. Solve for S=(2*2-1)/2=1.5 lb Ans Steel ball weighs 1.5 lbs, brass ball weighs 2lb